# Diverse, not Short: A Length-Controlled Self-Learning Framework for Improving Response Diversity of Language Models

### v2

@@ Abstract
Diverse language model responses are crucial for creative generation, open-ended tasks, and self-improvement training. We show that common diversity metrics, and even reward models used for preference optimization, systematically bias models toward shorter outputs, limiting expressiveness. To address this, we introduce Diverse, not Short (Diverse-NS), a length-controlled self-learning framework that improves response diversity while maintaining length parity. By generating and filtering preference data that balances diversity, quality, and length, Diverse-NS enables effective training using only 3,000 preference pairs. Applied to LLaMA-3.1-8B and the Olmo-2 family, Diverse-NS substantially enhances lexical and semantic diversity. We show consistent improvement in diversity with minor reduction or gains in response quality on four creative generation tasks: Divergent Associations, Persona Generation, Alternate Uses, and Creative Writing. Surprisingly, experiments with the Olmo-2 model family (7B, and 13B) show that smaller models like Olmo-2-7B can serve as effective “diversity teachers” for larger models. By explicitly addressing length bias, our method efficiently pushes models toward more diverse and expressive outputs.
Diverse, not Short: A Length-Controlled Self-Learning Framework for Improving Response Diversity of Language Models
@@ 1 Introduction
To address this overlooked confounding factor, we propose Diverse, not Short (Diverse-NS), a length-controlled self-learning framework that counteracts the hidden brevity bias in standard diversity metrics and improves diversity in both structured and free-form generation. The framework first uses sequential prompting to elicit more diverse responses, followed by preference pair curation that improve both diversity and quality while maintaining comparable response lengths (within words). Using these preference pairs, we apply Direct Preference Optimization (DPO) (Rafailov et al., 2023) to improve the response diversity of the base model. Our key contributions are:
@@
Diverse-NS: A length-controlled self-learning framework that significantly improves the response diversity of Llama-3.1-8B and Olmo-2-7B using only k preference pairs.
@@ 2 Related Work
Evaluation of diversity is challenging for a few reasons: length bias McCarthy and Jarvis (2010); Covington and McFall (2010); Mass (1972); Johnson et al. (2023), relative difficulty in achieving substantial agreement between humans (Chakrabarty et al., 2023, 2024; Gómez-Rodríguez and Williams, 2023), and inconsistent human preferences Evans et al. (2016). Despite the challenges, many studies have highlighted the compromised diversity of synthetic or human-LLM collaborative text Shaib et al. (2024b, a); Salkar et al. (2022); Padmakumar and He (2023); Guo et al. (2023); Kirk et al. (2023); Doshi and Hauser (2024); Anderson et al. (2024). So, we present a method, Diverse-NS, to increase the response diversity and propose a metric, Diversity Decile, to measure diversity in a length-controlled way.
@@ 4.1 Data Generation
We focus on a creative writing task to build the dataset for preference learning. The goal is to generate short stories (five sentences) that must include three words specified in the prompt. This task has been extensively validated in studies of human creativity (Prabhakaran et al., 2014). To create a diverse set of prompts, we first curated a list of 300 unique words, 111A manually curated list of 20 words was extended using GPT-4o and Claude-3.7.. For generating short stories from LMs, we create prompts by randomly sampling three-word sets from .
@@
Given the task setup, we create k story writing prompts, with k unique three-word sets. The exact prompt is provided in Section A.1. We initially sampled k stories (10 per prompt) using a temperature of from each of the following LMs: Llama-8B and Olmo-7B (Grattafiori et al., 2024; OLMo et al., 2024). Within the sampled stories, we extracted the repeating Part-Of-Speech (POS) bigrams and found that the start of the story is highly likely to have repetitions across different prompts (refer to Table B.1). To overcome these repetitions, we performed a second inference call to re-draft the story with additional constraints, an approach similar to Denial Prompting presented by Lu et al. (2024) (refer to Section A.1 exact prompt). In our case, unlike Lu et al. (2024), the constraints we use are specifically targeted to elicit a more diverse response from the model while maintaining the same (or comparable) length. With a pilot analysis on the initial k responses, we find that the story generated in the second inference call is on average more diverse (refer to Table B.2). These results motivated us to set up the final two-step data generation process, first inference call to collect natural responses from the model, and second inference call to redraft the natural response into a more diverse story. In the final data generation phase, we used k unique three-word sets to generate prompts and sampled 10 first and second responses for each prompt, resulting in a dataset of 200,000 tuples of prompt, first response, and second response, per model (Llama-8B and Olmo-7B). We denote the data as follows: where, , , and denote the prompt, first response, and second response, respectively, generated from model (policy) . Note that and we use two models, , for data generation.
@@ 4.2 Data Filtration
Each pair gives us a natural candidate for rejected and chosen responses. On average, the second response is more diverse than the first (Table B.2), but not every pair guarantees learning higher diversity. To ensure that the model receives consistent and useful learning signals, we apply a set of filtering rules.
@@
We use entropy to measure diversity and the ArmoRM reward model scores (Wang et al., 2024) to assess quality. Entropy is a standard metric for lexical diversity (Lanchantin et al., 2025), with higher values indicating greater diversity. In our self-learning setup, entropy is useful because it reflects the model’s likelihood of producing a certain continuation of the prompt. When used in filtering, it helps identify training data that aligns with the model’s own capabilities. For each example, we compute the entropy and the reward model score of both and , conditioned on the original prompt . When we use our data generation method, and use entropy and ArmoRM values for filtration, we call our approach, Diverse, not Short (Diverse-NS or D-NS).
@@
While entropy and ArmoRM scores are high-quality metrics for measuring diversity and response quality, they are computationally expensive. Each example requires two additional inference calls to compute entropy and two more for ArmoRM scoring. To reduce this overhead, we evaluated seven alternative metrics and measured their correlation with entropy and ArmoRM scores. Among these, Type-Token Ratio (TTR) showed the highest correlation with entropy (Pearson , ), and the MAAS index (Mass, 1972) was most correlated with ArmoRM scores (Pearson , ). Refer to Table 1 for all correlation results. Based on these findings, we replace entropy with TTR and ArmoRM scores with MAAS in our filtering pipeline. When this lightweight variant is used during data filtering, we refer to the resulting method as Diverse-NS-Lite (or D-NS-Lite).
@@
Based on the correlation analysis (Tab. 1), it is worth noting that both entropy and ArmoRM scores are negatively correlated with text length. As a result, optimizing for diversity or quality alone may unintentionally favor shorter responses as the “chosen” continuations. To avoid this bias, it is essential to explicitly control for length when curating preference learning data for improving diversity. To show this, we implement a recent study that is closest to our method, Diverse Preference Optimization(DivPO) (Lanchantin et al., 2025). DivPO also generates responses and filters the responses to form preference learning pairs without explicitly control the length of the chosen and rejected continuations. We compare pre- and post-filtration data properties for DivPO and Diverse-NS in Tab. 2. The table clearly shows that in the pursuit of maximizing the entropy values, DivPO selects significantly shorter (- words shorter on average) responses as the chosen responses in the final preference data.
@@ 5.1 Preference Tuning
After generating and filtering the data, we fine-tune the same base policy that was used to generate it. In other words, data generated by Llama-8B is used to train Llama-8B, and likewise for Olmo-7B. To ensure a fair comparison across methods (DivPO, D-NS, and D-NS-Lite), we limit the final training dataset to 3,000 preference pairs222We observed that the size of the dataset after filtering is the smallest for Diverse-NS, slightly more than k. Hence, to make the training runs more comparable across methods, we limit the size of the dataset to k for all methods.. To construct this k dataset, we first compute the entropy gain for each pair as the difference between the entropy of the chosen and rejected responses 333note that, by construction, the chosen response has higher entropy in the filtered set. We then sort all pairs by entropy gain in descending order and select the top k examples. This ensures that the final training set maximizes diversity gain for the base model. The same selection procedure is applied to all three methods.
We further extend our experiments to evaluate the utility of training larger models with data generated from smaller ones. For this, we train Olmo-13B using preference pairs generated from Olmo-7B. We provide all hyperparameter values in Appendix C.
All experiments are run on a single NVIDIA RTX 6000 GPU (48GB memory), using a per-device batch size of 2 and a global batch size of 64. Training Llama-8B or Olmo-7B takes approximately 100–150 minutes while O-13B takes 200-220 minutes per run, highlighting our setup efficiency.
@@ 5.2.1 Tasks
The DAT (Olson et al., 2021) is a psychological test commonly used to assess divergent thinking in humans. Participants are asked to generate a list of 10 words that are as dissimilar from each other as possible. Recent studies have adapted DAT to evaluate the creativity of language models, focusing on their ability to produce diverse outputs (Bellemare-Pepin et al., 2024). To quantify model performance on DAT, we use the Divergent Semantic Integration (DSI) metric (Johnson et al., 2023), which computes the average semantic distance of each word in the generated list from all others. Higher DSI values indicate more divergent thinking and greater ideological diversity. Following Johnson et al. (2023), we extract token embeddings from the layer of BERT-large for the generated list and compute the average pairwise cosine distance between all embeddings. This approach has been shown to correlate strongly with human judgments of creativity (Johnson et al., 2023). We provide the exact prompt used for DAT in Section A.2. For a robust evaluation, we sample 100 DAT responses per model using temperature 1.0 and different random seeds. From these 100 lists (each with 10 words), we compute and report two metrics: (1) the average and standard deviation of DSI scores, and (2) the number of unique words across all 1,000 generated tokens. In both cases, higher values indicate greater diversity.
@@
To assess diversity in structured generation, we use the PGT, also used in the study conducted by Lanchantin et al. (2025). In this task, the model is prompted to generate a JSON object with three fields: first name, city of birth, and current occupation to evaluate the model’s ability to produce varied persona descriptions. The exact prompt is provided in Section A.2. We sample 100 responses per model using temperature 1.0 and different random seeds. For each key in the JSON object, we report the proportion of unique values across the 100 responses. Higher uniqueness indicates greater diversity.
@@
The Alternate Uses Task (AUT) is a common and rigorously validated psychological test to measure human divergent thinking (Guilford, 1956). In this task, the subject/model is asked to generate creative and unconventional uses for objects (e.g., broom). The prompt and list of objects used for evaluation are provided in Section A.2. We use 15 unique objects and generate 10 responses per object using different random seeds, resulting in 150 total responses sampled at temperature 1.0. For quantifying the diversity of the generated uses, we measure the distance between the target object and generated uses with the help of BERT-large encodings, a validated approach that correlates with human creativity ratings (Patterson et al., 2023). We report the mean and standard deviation of the distance values, higher values indicate higher diversity.
@@
The CWT — based on a well-validated psychological assessment of creativity (Prabhakaran et al., 2014) — is exactly the same as our data generation task. That is, given a set of three words, the subject/model is tasked with generating a creative short story that includes all three words. We provide a separate list of three-word sets used for evaluation in Section A.2. We sample 10 responses for each of the seven three-word sets with temperature of . Unlike our other evaluation tasks, we measure the diversity as well as the quality of the generated responses. Similar to Johnson et al. (2023), we calculate the DSI metric to measure the diversity of the generated story. For quality measurements, we resort to the ArmoRM reward model preference scores (Wang et al., 2024). We report the average and standard deviation values of DSI and ArmoRM scores, and 4-gram diversity values, where higher values are more desirable for all metrics.
@@ 5.2.2 Length-Adjusted Evaluation
We first build a decile map that captures the empirical distribution of diversity scores at each length. Using 800 000 stories collected from Llama-8B and Olmo-7B over 40 000 prompts, we: (1) group responses by word count ; (2) compute decile thresholds for a chosen diversity metric (e.g. TTR, MTLD); and (3) store these percentile thresholds in a lookup table . Here, a decile refers to one of ten intervals that divide the distribution of diversity scores for a given length into ten equal parts. The top decile corresponds to the most diverse 10% of responses at that length, the second-highest to the next 10%, and so on. This mapping allows us to estimate the approximate diversity rank of any new response relative to other responses of the same length. At evaluation time, a new response with word count and diversity score is assigned the highest decile index such that exceeds the -th threshold in . Formally, where larger means the response is more diverse than a greater share of previously observed texts of the same length.
@@
Positive values indicate improved diversity, with higher values corresponding to a larger improvement. Negative values signify reduced diversity, and signifies no change. Note that, DD is agnostic to the choice of diversity metric. We therefore report DD values using seven standard metrics: TTR, MATTR, HD-D, MTLD, and MAAS. We also compute DD using ArmoRM reward scores to quantify the gain or loss in quality. This length-aware normalization prevents either long or short responses from being over-credited for diversity444We provide a summary of all metrics in Table G.1.

### v3

# Diverse, not Short: A Length-Controlled Data Selection Strategy for Improving Response Diversity of Language Models

@@ Abstract
Diverse language model responses are crucial for creative generation, open-ended tasks, and self-improvement training. We show that common diversity metrics, and even reward models used for preference optimization, systematically bias models toward shorter outputs, limiting expressiveness. To address this, we introduce Diverse, not Short (Diverse-NS), a length-controlled
data selection strategy
that improves response diversity while maintaining length parity. By generating and filtering preference data that balances diversity, quality, and length, Diverse-NS enables effective training using only 3,000 preference pairs. Applied to LLaMA-3.1-8B and the Olmo-2 family, Diverse-NS substantially enhances lexical and semantic diversity. We show consistent improvement in diversity with minor reduction or gains in response quality on four creative generation tasks: Divergent Associations, Persona Generation, Alternate Uses, and Creative Writing. Surprisingly, experiments with the Olmo-2 model family (7B, and 13B) show that smaller models like Olmo-2-7B can serve as effective “diversity teachers” for larger models. By explicitly addressing length bias, our method efficiently pushes models toward more diverse and expressive outputs111Code and data can be accessed at,
Code: Diverse-NS-Repo,
Dataset: Diverse-NS-Preference-Learning-Data.
Diverse, not Short: A Length-Controlled Data Selection Strategy for Improving Response Diversity of Language Models
@@ 1 Introduction
To address this overlooked confounding factor, we propose Diverse, not Short (Diverse-NS), a length-controlled data selection strategy that counteracts the hidden brevity bias in standard diversity metrics and improves diversity in both structured and free-form generation. The framework first uses sequential prompting to elicit more diverse responses, followed by preference pair curation that improve both diversity and quality while maintaining comparable response lengths (within words). Using these preference pairs, we apply Direct Preference Optimization (DPO) (Rafailov et al., 2023) to improve the response diversity of the base model. Our key contributions are:
@@
Diverse-NS: A length-controlled data selection strategy that significantly improves the response diversity of Llama-3.1-8B and Olmo-2-7B using only k preference pairs.
@@ 2 Related Work
Evaluation of diversity is challenging for a few reasons: length bias McCarthy and Jarvis (2010); Covington and McFall (2010); Mass (1972); Johnson et al. (2023); Deshpande et al. (2025), relative difficulty in achieving substantial agreement between humans (Chakrabarty et al., 2023, 2024; Gómez-Rodríguez and Williams, 2023), and inconsistent human preferences Evans et al. (2016). Despite the challenges, many studies have highlighted the compromised diversity of synthetic or human-LLM collaborative text Shaib et al. (2024b, a); Salkar et al. (2022); Padmakumar and He (2023); Guo et al. (2023); Kirk et al. (2023); Doshi and Hauser (2024); Anderson et al. (2024). So, we present Diverse-NS, to increase the response diversity and propose a metric, Diversity Decile, to measure diversity in a length-controlled way.
@@ 4.1 Data Generation
We focus on a creative writing task to build the dataset for preference learning. The goal is to generate short stories (five sentences) that must include three words specified in the prompt. This task has been extensively validated in studies of human creativity (Prabhakaran et al., 2014). To create a diverse set of prompts, we first curated a list of 300 unique words, 222A manually curated list of 20 words was extended using GPT-4o and Claude-3.7.. For generating short stories from LMs, we create prompts by randomly sampling three-word sets from .
@@
Given the task setup, we create k story writing prompts, with k unique three-word sets. The exact prompt is provided in Section˜A.1. We initially sampled k stories (10 per prompt) using a temperature of from each of the following LMs: Llama-8B and Olmo-7B (Grattafiori et al., 2024; OLMo et al., 2024). Within the sampled stories, we extracted the repeating Part-Of-Speech (POS) bigrams and found that the start of the story is highly likely to have repetitions across different prompts (refer to Table˜B.1). To overcome these repetitions, we performed a second inference call to re-draft the story with additional constraints, an approach similar to Denial Prompting presented by Lu et al. (2024) (refer to Section˜A.1 exact prompt). In our case, unlike Lu et al. (2024), the constraints we use are specifically targeted to elicit a more diverse response from the model while maintaining the same (or comparable) length. With a pilot analysis on the initial k responses, we find that the story generated in the second inference call is on average more diverse (refer to Table˜B.2). These results motivated us to set up the final two-step data generation process, first inference call to collect natural responses from the model, and second inference call to redraft the natural response into a more diverse story. In the final data generation phase, we used k unique three-word sets to generate prompts and sampled 10 first and second responses for each prompt, resulting in a dataset of 200,000 tuples of prompt, first response, and second response, per model (Llama-8B and Olmo-7B). We denote the data as follows: where, , , and denote the prompt, first response, and second response, respectively, generated from model (policy) . Note that and we use two models, , for data generation.
@@ 4.2 Data Filtration
Each pair gives us a natural candidate for rejected and chosen responses. On average, the second response is more diverse than the first (Table˜B.2), but not every pair guarantees learning higher diversity. To ensure that the model receives consistent and useful learning signals, we apply a set of filtering rules.
@@
We use entropy to measure diversity and the ArmoRM reward model scores (Wang et al., 2024) to assess quality333refer to Appendix C for a brief description of ArmoRM scores. Entropy is a standard metric for lexical diversity (Lanchantin et al., 2025), with higher values indicating greater diversity. In our self-learning setup, entropy is useful because it reflects the model’s likelihood of producing a certain continuation of the prompt. When used in filtering, it helps identify training data that aligns with the model’s own capabilities. For each example, we compute the entropy and the reward model score of both and , conditioned on the original prompt . When we use our data generation method, and use entropy and ArmoRM values for filtration, we call our approach, Diverse, not Short (Diverse-NS or D-NS).
@@
While entropy and ArmoRM scores are high-quality metrics for measuring diversity and response quality, they are computationally expensive. Each example requires two additional inference calls to compute entropy and two more for ArmoRM scoring. To reduce this overhead, we evaluated seven alternative metrics and measured their correlation with entropy and ArmoRM scores. Among these, Type-Token Ratio (TTR) showed the highest correlation with entropy (Pearson , ), and the MAAS index (Mass, 1972) was most correlated with ArmoRM scores (Pearson , ). Refer to Table˜1 for all correlation results. Based on these findings, we replace entropy with TTR and ArmoRM scores with MAAS in our filtering pipeline. When this lightweight variant is used during data filtering, we refer to the resulting method as Diverse-NS-Lite (or D-NS-Lite). We provide a discussion on the computational savings achieved with Diverse-NS-Lite in Appendix˜D.
@@
Based on the correlation analysis (Table˜1), it is worth noting that both entropy and ArmoRM scores are negatively correlated with text length. As a result, optimizing for diversity or quality alone may unintentionally favor shorter responses as the “chosen” continuations. To avoid this bias, it is essential to explicitly control for length when curating preference learning data for improving diversity. To show this, we implement a recent study that is closest to our method, Diverse Preference Optimization(DivPO) (Lanchantin et al., 2025) (refer to Appendix˜E for implementation details). DivPO also generates responses and filters the responses to form preference learning pairs without explicitly controlling the length of the chosen and rejected continuations. We compare pre- and post-filtration data properties for DivPO and Diverse-NS in Tab. 2. The table clearly shows that in the pursuit of maximizing the entropy values, DivPO selects significantly shorter (- words shorter on average) responses as the chosen responses in the final preference data.
@@ 5.1 Preference Tuning
After generating and filtering the data, we fine-tune the same base policy that was used to generate it. In other words, data generated by Llama-8B is used to train Llama-8B, and likewise for Olmo-7B. To ensure a fair comparison across methods (DivPO, D-NS, and D-NS-Lite), we limit the final training dataset to 3,000 preference pairs444We observed that the size of the dataset after filtering is the smallest for Diverse-NS, slightly more than k. Hence, to make the training runs more comparable across methods, we limit the size of the dataset to k for all methods.. To construct this k dataset, we first compute the entropy gain for each pair as the difference between the entropy of the chosen and rejected responses 555note that, by construction, the chosen response has higher entropy in the filtered set. We then sort all pairs by entropy gain in descending order and select the top k examples. This ensures that the final training set maximizes diversity gain for the base model. The same selection procedure is applied to all three methods.
We further extend our experiments to evaluate the utility of training larger models with data generated from smaller ones. For this, we train Olmo-13B using preference pairs generated from Olmo-7B. We provide all hyperparameter values in Appendix˜F. All experiments are run on a single NVIDIA RTX 6000 GPU (48GB memory), using a per-device batch size of 2 and a global batch size of 64. Training Llama-8B or Olmo-7B takes approximately 100–150 minutes while O-13B takes 200-220 minutes per run, highlighting our setup efficiency.
@@ 5.2.1 Tasks
The DAT (Olson et al., 2021) is a psychological test commonly used to assess divergent thinking in humans. Participants are asked to generate a list of 10 words that are as dissimilar from each other as possible. Recent studies have adapted DAT to evaluate the creativity of language models, focusing on their ability to produce diverse outputs (Bellemare-Pepin et al., 2024). To quantify model performance on DAT, we use the Divergent Semantic Integration (DSI) metric (Johnson et al., 2023), which computes the average semantic distance of each word in the generated list from all others. Higher DSI values indicate more divergent thinking and greater ideological diversity. Following Johnson et al. (2023), we extract token embeddings from the layer of BERT-large for the generated list and compute the average pairwise cosine distance between all embeddings. This approach has been shown to correlate strongly with human judgments of creativity (Johnson et al., 2023). We provide the exact prompt used for DAT in Section˜A.2. For a robust evaluation, we sample 100 DAT responses per model using temperature 1.0 and different random seeds. From these 100 lists (each with 10 words), we compute and report two metrics: (1) the average and standard deviation of DSI scores, and (2) the number of unique words across all 1,000 generated tokens. In both cases, higher values indicate greater diversity.
@@
To assess diversity in structured generation, we use the PGT, also used in the study conducted by Lanchantin et al. (2025). In this task, the model is prompted to generate a JSON object with three fields: first name, city of birth, and current occupation to evaluate the model’s ability to produce varied persona descriptions. The exact prompt is provided in Section˜A.2. We sample 100 responses per model using temperature 1.0 and different random seeds. For each key in the JSON object, we report the proportion of unique values across the 100 responses. Higher uniqueness indicates greater diversity.
@@
The Alternate Uses Task (AUT) is a common and rigorously validated psychological test to measure human divergent thinking (Guilford, 1956). In this task, the subject/model is asked to generate creative and unconventional uses for objects (e.g., broom). The prompt and list of objects used for evaluation are provided in Section˜A.2. We use 15 unique objects and generate 10 responses per object using different random seeds, resulting in 150 total responses sampled at temperature 1.0. For quantifying the diversity of the generated uses, we measure the distance between the target object and generated uses with the help of BERT-large encodings, a validated approach that correlates with human creativity ratings (Patterson et al., 2023). We report the mean and standard deviation of the distance values, higher values indicate higher diversity.
@@
The CWT — based on a well-validated psychological assessment of creativity (Prabhakaran et al., 2014) — is exactly the same as our data generation task. That is, given a set of three words, the subject/model is tasked with generating a creative short story that includes all three words. We provide a separate list of three-word sets used for evaluation in Section˜A.2. We sample 10 responses for each of the seven three-word sets with temperature of . Unlike our other evaluation tasks, we measure the diversity as well as the quality of the generated responses. Similar to Johnson et al. (2023), we calculate the DSI metric to measure the diversity of the generated story. For quality measurements, we resort to the ArmoRM reward model preference scores (Wang et al., 2024). We report DSI, ArmoRM, and 4-gram diversity values, where higher values are more desirable for all metrics.
@@ 5.2.2 Length-Adjusted Evaluation
We first build a decile map that captures the empirical distribution of diversity scores at each length. Using 800 000 stories collected from Llama-8B and Olmo-7B over 40 000 prompts, we: (1) group responses by word count ; (2) compute decile thresholds for a chosen diversity metric (e.g. TTR, MTLD); and (3) store these percentile thresholds in a lookup table . Here, a decile refers to one of ten intervals that divide the distribution of diversity scores for a given length into ten equal parts. The top decile corresponds to the most diverse 10% of responses at that length, the second-highest to the next 10%, and so on. This mapping allows us to estimate the approximate diversity rank of any new response relative to other responses of the same length. At evaluation time, a new response with word count and diversity score is assigned the highest decile index such that exceeds the -th threshold in . Formally, where larger means the response is more diverse than a greater share of previously observed texts of the same length666If the response length of the new response is not present in the constructed map, we select the closest neighbor as an approximation. However, it is important to note that due to the ample size of our synthetic data, we never encounter a case where the response length value is not found in the constructed map..
@@
Positive values indicate improved diversity, with higher values corresponding to a larger improvement. Negative values signify reduced diversity, and signifies no change. Note that, DD is agnostic to the choice of diversity metric. We therefore report DD values using seven standard metrics: TTR, MATTR, HD-D, MTLD, and MAAS. We also compute DD using ArmoRM reward scores to quantify the gain or loss in quality. This length-aware normalization prevents either long or short responses from being over-credited for diversity777We provide a summary of all metrics in Table J.1.